# An Effective Deligne’s equidistribution theorem

### v1

@@ 1. Introduction
Let $\mathbf{F}_{q}$ be a finite field of characteristic $p$ with $q$ elements, $X$ a smooth geometrically connected algebraic curve over $\mathbf{F}_{q}$, $\eta$ the generic point of $X$, $\ell$ a prime number distinct from $p$, and $\mathcal{F}$ a lisse $\overline{\mathbf{Q}}_{\ell}$-sheaf on $X$. We have an exact sequence of fundamental groups
@@
We assume $G$ is connected throughout this paper. Let $\mathfrak{g}$ be the Lie algebra of $G_{\mathbf{C}}$. Fix a Cartan subalgebra $\mathfrak{h}$ of $\mathfrak{g}$, and let $R$ be the root system. Choose an ordering $R=R^{+}\cup R^{-}$ on $R$. For each ${\boldsymbol{\alpha}}\in R$, let $H_{\boldsymbol{\alpha}}\in\mathfrak{h}$ be the element in $[\mathfrak{g}_{\boldsymbol{\alpha}},\mathfrak{g}_{-{\boldsymbol{\alpha}}}]$ satisfying ${\boldsymbol{\alpha}}(H_{\boldsymbol{\alpha}})=2$, and let $\mathfrak{h}_{\mathbf{R}}$ be the real subspace of $\mathfrak{h}$ spanned by $H_{\boldsymbol{\alpha}}$ $(\boldsymbol{\alpha}\in R)$. There exists a real subalgebra $\mathfrak{g}_{0}$ of $\mathfrak{g}$ such that
$\mathfrak{g}_{0}=i\mathfrak{h}_{\mathbf{R}}\bigoplus\bigoplus_{{\boldsymbol{\alpha}}\in R^{+}}\Big{(}\mathfrak{g}_{0}\cap(\mathfrak{g}_{\boldsymbol{\alpha}}\oplus\mathfrak{g}_{-{\boldsymbol{\alpha}}})\Big{)},$
that $\mathfrak{g}_{0}$ is the Lie algebra of a maximal compact subgroup ${G_{\mathbf{R}}}$ of $G$, and that $i\mathfrak{h}_{\mathbf{R}}$ is the Lie algebra of a maximal torus $T$ of ${G_{\mathbf{R}}}$. The category of finite dimensional complex representations of $G_{\mathbf{C}}$ is equivalent to that of finite dimensional complex representations of $G_{\mathbf{R}}$.
@@
|$\frac{|\{x\in X(\mathbf{F}_{q^{m}}):\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}}\in D\}|}{|X(\mathbf{F}_{q^{m}})|}=\mu_{G_{\mathbf{R}}^{\natural}}(D)+O\Big{(}q^{-\frac{m}{2(|R^{+}|+n)}}\Big{)},$
where $n=\mathrm{dim}_{\mathbf{C}}\mathfrak{h}$ is the rank of the Lie group, and the constant implied by $O$ depends only on $X$ and $\mathcal{F}$.
We will describe small boxes in Definition 1.3 below. We deduce the above theorem from an Erdős-Turán type inequality for compact Lie groups, which is of independent interest and we now describe. Let
@@
Hence $W$ acts on the Lie algebra $\mathfrak{h}_{\mathbf{R}}$ and its dual $\mathfrak{h}_{\mathbf{R}}^{*}$. Provide $\mathfrak{h}_{\mathbf{R}}$ and $\mathfrak{h}_{\mathbf{R}}^{*}$ with the inner product induced by the Killing form. The Weyl group can be identified with the subgroup of $\mathrm{GL}(\mathfrak{h}_{\mathbf{R}})$ generated by reflections with respect to $H_{\boldsymbol{\alpha}}^{\perp}$ $({\boldsymbol{\alpha}}\in R)$, and with the subgroup of $\mathrm{GL}(\mathfrak{h}_{\mathbf{R}}^{*})$ generated by reflections with respect to ${\boldsymbol{\alpha}}^{\perp}$ $({\boldsymbol{\alpha}}\in R)$. Let $T/W$ be the quotient space. The map $T\to G_{\mathbf{R}}^{\natural}$ sending each element in $T$ to its conjugacy class in $G_{\mathbf{R}}$ defines a homeomorphism
@@
Any representation $G\to\mathrm{GL}(V)$ gives rise to a representation of $\mathfrak{g}\to\mathfrak{gl}(V)$ such that the following diagram commutes:
$\begin{array}[]{ccccc}2\pi i\mathfrak{h}_{\mathbf{R}}&\hookrightarrow&\mathfrak{g}_{0}&\to&\mathfrak{gl}(V)\\ {\scriptstyle\exp}\downarrow&&{\scriptstyle\exp}\downarrow&&{\scriptstyle\exp}\downarrow\\ T&\hookrightarrow&G_{0}&\to&\mathrm{GL}(V).\end{array}$
Let ${\boldsymbol{\lambda}}\in\mathfrak{h}^{*}$ be a weight of the representation. Then there exists a nonzero vector $v\in V$ such that $Hv={\boldsymbol{\lambda}}(H)v$ for all $H\in\mathfrak{h}$. We thus have
$\exp(2\pi iH)(v)=e^{2\pi i{\boldsymbol{\lambda}}(H)}v.$
The map $\exp:2\pi i\mathfrak{h}_{\mathbf{R}}\to T$ is an epimormphism. Let $2\pi i\Gamma$ be its kernel. We call $\Gamma$ the integral lattice. If $H\in\Gamma$, then we have $e^{2\pi i{\boldsymbol{\lambda}}(H)}v=v$, and hence $e^{2\pi i{\boldsymbol{\lambda}}(H)}=1$. So we have
${\boldsymbol{\lambda}}(\Gamma)\subset\mathbf{Z}.$
Define the lattice of integral forms $I^{*}$ to be
@@
In particular, for any root ${\boldsymbol{\alpha}}\in R$, we have a character $e({\boldsymbol{\alpha}})$ of $T$. Define the weight lattice to be
@@
$E=\Big{\{}\exp\Big{(}2\pi i\sum_{j=1}^{n}t_{j}\mathbf{e}_{j}\Big{)}:\,t_{j}\in I_{j}\Big{\}}$
for some intervals $I_{j}$ of lengths $<1$ is called a box in $T$. We say the box $E$ is small if for any nonidentity element $\sigma$ in the Weyl group $W$, $E\cap\sigma(E)$ is empty. A subset $D$ of $T/W$ (resp. $G_{\mathbf{R}}^{\natural}$) is called a small box if it is the image of a small box $E$ in $T$ under the map
@@
With the above notations, since the lengths of the intervals $I_{j}$ are $<1$, the map $\exp(2\pi i\hbox{-}):\mathfrak{h}_{R}\to T$ identifies the set $I_{1}\times\cdots\times I_{n}$ with its image $E$ in $T$. Suppose furthermore that $E\cap\sigma(E)$ is empty for every nonidentity element $\sigma\in W$. Then the projection $T\to T/W$ identifies $E$ with its image $D$ in $T/W$, and the inverse image of $D$ in $T$ is $\bigsqcup_{\sigma\in W}\sigma(E)$.
@@
where for any $\sigma\in W$, $\mathrm{sgn}(\sigma)$ is defined to be determinant of $\sigma$ considered as a linear operator on $\mathfrak{h}_{\mathbf{R}}^{*}$. Its values are ±$\pm 1$ since $\sigma$ is a composite of reflections. Let
@@
Let $\boldsymbol{\boldsymbol{\alpha}}_{1},\ldots,\boldsymbol{\boldsymbol{\alpha}}_{n}\in R^{+}$ be all the simple roots. The fundamental weights $\boldsymbol{\omega}_{1},\ldots,\boldsymbol{\omega}_{n}$ in the weight lattice $\Lambda$ are defined by
$\boldsymbol{\omega}_{i}(H_{\boldsymbol{\boldsymbol{\alpha}}_{j}})=\delta_{ij}.$
Any $\boldsymbol{\boldsymbol{\lambda}}\in\mathfrak{h}^{*}_{\mathbf{R}}$ can be written as $\boldsymbol{\boldsymbol{\lambda}}=\sum_{j=1}^{n}{\lambda}_{j}\boldsymbol{\omega}_{j}$ with $\lambda_{j}\in\mathbf{R}$, and $\boldsymbol{\boldsymbol{\lambda}}$ lies in $\Lambda\cap\mathfrak{C}$ if and only if ${\lambda}_{j}$ are nonnegative integers. The following is a generation of the Erdős–Turán inequality ([Mo94, Chapter 1, Corollary 1.1]) to compact Lie groups.
@@
$\ll$ $\frac{1}{N}\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}-\{0\}\\ 0\leq\lambda_{j}\leq M\end{subarray}}\frac{1}{\|\boldsymbol{\lambda}\|+1}\Big{|}\sum_{i=1}^{N}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}(x_{i})\Big{|}+\frac{1}{M},$
where $\|\boldsymbol{\lambda}\|=\sqrt{\sum_{i=1}^{n}\lambda_{i}^{2}},$ and the constant implied by $\ll$ depends only on $G$.
@@
For $G=\mathrm{SL}_{2}$, Theorem 1.5 is due to Niederreiter [Ni91, Lemma 3]. Rouse and Thorner [RT17, Lemma 3.1] give another proof of Niederreiter’s theorem using the Beurling and Selberg majorizing and minorizing functions. We prove Theorem 1.5 using Cochrane’s $n$-dimensional version of the Beurling and Selberg functions ([Co85]).
@@
$=$ $\mu_{G^{\natural}_{\mathbf{R}}}(D)\mu_{G^{\prime\natural}_{\mathbf{R}}}(D^{\prime})+O\Big{(}q^{-\frac{m}{2(|R^{+}|+|R^{\prime+}|+n+n^{\prime}-1)}}\Big{)},$

### v2

@@ 1. Introduction
Let $\mathbf{F}_{q}$ be a finite field of characteristic $p$ with $q$ elements, $X$ a smooth geometrically connected scheme over $\mathbf{F}_{q}$, $\eta$ the generic point of $X$, $\ell$ a prime number distinct from $p$, and $\mathcal{F}$ a lisse $\overline{\mathbf{Q}}_{\ell}$-sheaf on $X$. We have an exact sequence of fundamental groups
@@
We assume $G$ is connected throughout this paper. Let $\mathfrak{g}$ be the Lie algebra of $G_{\mathbf{C}}$. Fix a Cartan subalgebra $\mathfrak{h}$ of $\mathfrak{g}$, and let $R$ be the root system. Choose an ordering $R=R^{+}\cup R^{-}$ on $R$. For each ${\boldsymbol{\alpha}}\in R$, let $H_{\boldsymbol{\alpha}}\in\mathfrak{h}$ be the element in $[\mathfrak{g}_{\boldsymbol{\alpha}},\mathfrak{g}_{-{\boldsymbol{\alpha}}}]$ satisfying ${\boldsymbol{\alpha}}(H_{\boldsymbol{\alpha}})=2$, and let $\mathfrak{h}_{\mathbf{R}}$ be the real subspace of $\mathfrak{h}$ spanned by $H_{\boldsymbol{\alpha}}$ $(\boldsymbol{\alpha}\in R)$. There exists a real subalgebra $\mathfrak{g}_{\mathbf{R}}$ of $\mathfrak{g}$ such that
$\mathfrak{g}_{\mathbf{R}}=i\mathfrak{h}_{\mathbf{R}}\bigoplus\bigoplus_{{\boldsymbol{\alpha}}\in R^{+}}\Big{(}\mathfrak{g}_{\mathbf{R}}\cap(\mathfrak{g}_{\boldsymbol{\alpha}}\oplus\mathfrak{g}_{-{\boldsymbol{\alpha}}})\Big{)},$
that $\mathfrak{g}_{\mathbf{R}}$ is the Lie algebra of a maximal compact subgroup ${G_{\mathbf{R}}}$ of $G$, and that $i\mathfrak{h}_{\mathbf{R}}$ is the Lie algebra of a maximal torus $T$ of ${G_{\mathbf{R}}}$. The category of finite dimensional complex representations of $G_{\mathbf{C}}$ is equivalent to that of finite dimensional complex representations of $G_{\mathbf{R}}$.
@@
|$\frac{|\{x\in X(\mathbf{F}_{q^{m}}):\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}}\in D\}|}{|X(\mathbf{F}_{q^{m}})|}=\mu_{G_{\mathbf{R}}^{\natural}}(D)+O\Big{(}q^{-\frac{m}{2(|R^{+}|+1)}}(\log q^{m})^{\frac{n-1}{|R^{+}|+1}}\Big{)},$
where $n:=\mathrm{dim}_{\mathbf{C}}\mathfrak{h}$ is the rank of the Lie group, and the constant implied by $O$ depends only on $X$ and $\mathcal{F}$.
We will describe small boxes in Definition 1.3 below. In the case where $X$ is an algebraic curve, let $\overline{X}$ be a smooth compactification of $X$. Then the constant implied by $O$ in the theorem depends only on the group $G$, the genus of $\overline{X}$, the number of points in $(\overline{X}-X)(\overline{\mathbf{F}}_{q})$, the rank of $\mathcal{F}$ and the Swan conductors of $\mathcal{F}|_{I_{x}}$ ($x\in\overline{X}-X$), where $I_{x}$ is the inertia subgroup of $\pi_{1}(X,\bar{\eta})$ at $x$.
We deduce the above theorem from an Erdős-Turán type inequality for compact Lie groups, which is of independent interest and we now describe. Let
@@
Hence $W$ acts on the Lie algebra $\mathfrak{h}_{\mathbf{R}}$ and its dual $\mathfrak{h}_{\mathbf{R}}^{*}$. Provide $\mathfrak{h}_{\mathbf{R}}$ and $\mathfrak{h}_{\mathbf{R}}^{*}$ with the inner product induced by the Killing form. The Weyl group can be identified with the subgroup of $\mathrm{GL}(\mathfrak{h}_{\mathbf{R}}^{*})$ generated by reflections with respect to ${\boldsymbol{\alpha}}^{\perp}$ $({\boldsymbol{\alpha}}\in R)$. Let $T/W$ be the quotient space. The map $T\to G_{\mathbf{R}}^{\natural}$ sending each element in $T$ to its conjugacy class in $G_{\mathbf{R}}$ defines a homeomorphism
@@
The map $\exp:2\pi i\mathfrak{h}_{\mathbf{R}}\to T$ is an epimormphism. Let $2\pi i\Gamma$ be its kernel. Then we have an isomorphism
$\exp(2\pi i\hbox{-}):\mathfrak{h}_{\mathbf{R}}/\Gamma\stackrel{{\scriptstyle\cong}}{{\to}}T.$
We call $\Gamma$ the integral lattice. Define the lattice of integral forms $I^{*}$ to be
@@
For any weight ${\boldsymbol{\lambda}}\in\mathfrak{h}^{*}$ of a representation of $\mathfrak{g}$ coming from a representation of $G$, we have $\boldsymbol{\lambda}\in I^{*}$. In particular, for any root ${\boldsymbol{\alpha}}\in R$, we have $\alpha\in I^{*}$ and we have a character $e({\boldsymbol{\alpha}})$ of $T$. Define the weight lattice to be
@@
$\Omega=\Big{\{}\exp\Big{(}2\pi i\sum_{j=1}^{n}t_{j}\mathbf{e}_{j}\Big{)}:\,t_{j}\in I_{j}\Big{\}}$
for some intervals $I_{j}$ of lengths $<1$ is called a box in $T$. We say the box $\Omega$ is small if for any nonidentity element $\sigma$ in the Weyl group $W$, $\Omega\cap\sigma(\Omega)$ is empty. A subset $D$ of $T/W$ (resp. $G_{\mathbf{R}}^{\natural}$) is called a small box if it is the image of a small box $\Omega$ in $T$ under the canonical map
@@
With the above notations, since the lengths of the intervals $I_{j}$ are $<1$, the map $\exp(2\pi i\hbox{-}):\mathfrak{h}_{R}\to T$ identifies the set $I_{1}\times\cdots\times I_{n}$ with its image $\Omega$ in $T$. Suppose furthermore that $\Omega\cap\sigma(\Omega)$ is empty for every nonidentity element $\sigma\in W$. Then the projection $T\to T/W$ identifies $\Omega$ with its image $D$ in $T/W$, and the inverse image of $D$ in $T$ is $\bigsqcup_{\sigma\in W}\sigma(\Omega)$.
@@
where for any $\sigma\in W$, $\mathrm{sgn}(\sigma)$ is defined to be determinant of $\sigma$ considered as a linear operator on $\mathfrak{h}_{\mathbf{R}}^{*}$. Its value is $1$ or $-1$ since $\sigma$ is a composite of reflections. Let
@@
The basis $\{\mathbf{e}_{1},\ldots,\mathbf{e}_{n}\}$ of $\Gamma$ in Definition 1.3 is also a basis of $\mathfrak{h}_{\mathbf{R}}$, which we fix henceforth. Let $\{\mathbf{e}_{1}^{*},\ldots,\mathbf{e}_{n}^{*}\}$ be the dual basis for $\mathfrak{h}^{*}_{\mathbf{R}}$. For any $\boldsymbol{\lambda}\in\mathfrak{h}^{*}_{\mathbf{R}}$, write $\boldsymbol{\boldsymbol{\lambda}}=\sum_{j=1}^{n}{\lambda}_{j}\mathbf{e}^{*}_{j}$ and let
(1.4.1) ${\mathbf{N}}(\boldsymbol{\lambda}):=(|\lambda_{1}|+1)\cdots(|\lambda_{n}|+1).$
Finally, fix a norm
(1.4.2) $\|\cdot\|:\mathfrak{h}_{\mathbf{R}}\to\mathbf{R}_{\geq 0}$
on $\mathfrak{h}_{\mathbf{R}}.$ The following is a generation of the Erdős–Turán inequality ([Mo94, Chapter 1, Corollary 1.1]) to compact Lie groups.
@@
$\ll$ $\frac{1}{N}\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}-\{\mathbf{0}\}\\ \|\boldsymbol{\lambda}\|\leq M\end{subarray}}c(\boldsymbol{\lambda})\Big{|}\sum_{i=1}^{N}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}(x_{i})\Big{|}+\frac{1}{M},$
where $c(\boldsymbol{\lambda})=\sum_{\sigma\in W}\frac{1}{{\mathbf{N}}(\sigma(\boldsymbol{\lambda}))}$. For any real number $r$, we have
(1.5.1) $\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}-\{\mathbf{0}\}\\ \|\boldsymbol{\lambda}\|\leq M\end{subarray}}c(\boldsymbol{\lambda})\|\boldsymbol{\lambda}\|^{r}\ll M^{r}(\log M)^{n-1}.$
The constants implied by $\ll$ depend only on $G$.
@@
For $G=\mathrm{SL}_{2}$, Theorem 1.5 is due to Niederreiter [Ni91, Lemma 3]. Rouse and Thorner [RT17, Lemma 3.1] give another proof of Niederreiter’s theorem using the Beurling and Selberg majorizing and minorizing functions. We prove Theorem 1.5 using the general theory of the majorizing and minorizing functions of Colzani-Gigante-Travaglini ([CGT11]). In [R13], Rosengarten obtains a version of the Erdős–Turán inequality for simply connected compact Lie groups which is different from ours.
@@
$=\mu_{G^{\natural}_{\mathbf{R}}}(D)\mu_{G^{\prime\natural}_{\mathbf{R}}}(D^{\prime})+O\Big{(}q^{-\frac{m}{2(|R^{+}|+|R^{\prime+}|+1)}}(\log q^{m})^{\frac{n+n^{\prime}-2}{|R^{+}|+|R^{\prime+}|+1}}\Big{)},$